What are the roots of the quadratic equation x² + 4x + 7 = 0?

Grade:7

2 Answers

Apptica Corporations
10 years ago
This is a equation with imaginary roots. Discriminant<0 -2+i root3, -2-i root 3
bharat bajaj IIT Delhi
10 years ago
Discriminant = b^2 - 4ac = 4^2 - 4(1)(7) = 16 - 28 = -12 < 0
As D<0, the equation has complex roots.

x = -4 + v12 i / 2 = -2 + v3i
x = -2 - v3i
These are the roots of this equation where i is iota.
Thanks
Bharat Bajaj
IIT Delhi
askiitians faculty